# Sensitivity Analysis for Climate Science with Generative Flow Models

### v1

@@ Appendix B Getting full derivatives with the adjoint sensitivity method
from $t=0$ to $t=T$, where $w_{T}=\frac{\partial q}{\partial c}$ and $v_{T}=\frac{\partial q}{\partial\tau}$. In practice, calculating the additional derivative $\frac{\partial u_{\theta}}{\partial\tau}$ adds negligible computational cost, as the backpropagation process through the computational graph happens the same way and with the same starting tangent $a_{t}$. The total derivative is then
@@
In the case of cBottle, we know that $c(\tau)$ is linearly interpolated between the nearest two mid-month monthly mean SSTs (see Appendix C.2)
and so
$\frac{\partial\tau}{\partial c}=\left(\frac{\partial c}{\partial\tau}\right)^{-1}=\frac{\tau_{i+1}-\tau_{i}}{c(\tau_{i+1})-c(\tau_{i})}.$
Otherwise, a more general method of approximating $\frac{\partial\tau}{\partial c}$ is to train a regressor and to get this gradient with an AD call.
@@ C.2 The nature of SST conditioning data
As mentioned in the results section, cBottle is using one of its conditioning variables, the day of year $\tau$, to obtain the sea surface temperature forcing $c(\tau)$. This procedure is mentioned but not detailed in their paper, and so we will describe here. They use the AMIP dataset input4MIPs.CMIP6Plus.CMIP.PCMDI.PCMDI-AMIP-1-1-9.ocean.mon.tosbcs.g which can be downloaded with a script from here. This dataset contains the estimated global mid-month, monthly mean SST from January 1870 to December 2022, all dated on the 16th of the month (except February, which has the 15th). If the model is given a day of year $\tau$, $c(\tau)$ is linearly interpolated from the nearest two data points:
@@
Figure 5: A 3D slice of the SST manifold used to train cBottle, the years 1991 to 2021 are shown. Because of the high density of interpolated values (here denoted by thin grey lines), we expect gradients obtained in an SST change direction corresponding to seasonal SST changes to be sensible. The same cannot be said for directions pointing away from the manifold, which most of them will be.

### v2

@@ Appendix A Getting total derivatives with the adjoint sensitivity method
from $t=0$ to $t=T$, where $w_{T}=\frac{\partial q}{\partial c}$ and $v_{T}=\frac{\partial q}{\partial\tau}$. In practice, calculating the additional derivative $\frac{\partial u_{\theta}}{\partial\tau}$ adds negligible computational cost, as the backpropagation process through the computational graph happens the same way and with the same starting tangent $a_{t}$. Any input of the model can be added to the ODE this way. The total derivative is then
@@
To approximate $\frac{\partial\tau}{\partial c}$, one could train a regressor and get this gradient with an AD call.
For calculating the RMSE on historical data this is not needed, as we have both $\delta c$ and $\delta\tau$:
$\frac{\text{d}q}{\text{d}c}\cdot\delta c=\frac{\partial q}{\partial c}\cdot\delta c+\frac{\partial q}{\partial\tau}\frac{\partial\tau}{\partial c}\cdot\delta c\approx w_{T}\cdot\delta c+v_{T}\delta\tau.$
Appendix B More sensitivity maps
See Figure 3 for monthly sensitivity maps, each one averaged over $K/12\approx 211$ samples.
Figure 3: Monthly averaged sensitivity maps $\frac{\partial\bar{N}}{\partial\text{SST}}$ in the period 1971-2020.
@@ C.2 The nature of SST conditioning data
Climate in a Bottle is using one of its conditioning variables, the day of year $\tau$, to obtain the sea surface temperature forcing $c(\tau)$. This procedure is mentioned but not detailed in their paper, and so we will describe here. They use the AMIP dataset input4MIPs.CMIP6Plus.CMIP.PCMDI.PCMDI-AMIP-1-1-9.ocean.mon.tosbcs.g which can be downloaded with a script from here. This dataset contains the estimated global mid-month, monthly mean SST from January 1870 to December 2022, all dated on the 16th of the month (except February, which has the 15th). If the model is given a day of year $\tau$, $c(\tau)$ is linearly interpolated from the nearest two data points: